# Analysis of Pseudo-Labeling for Online Source-Free Universal Domain Adaptation

### v1

@@ Abstract
A domain (distribution) shift between training and test data often hinders the real-world performance of deep neural networks, necessitating unsupervised domain adaptation (UDA) to bridge this gap. Online source-free UDA has emerged as a solution for practical scenarios where access to source data is restricted and target data is received as a continuous stream. However, the open-world nature of many real-world applications additionally introduces category shifts meaning that the source and target label spaces may differ. Online source-free universal domain adaptation (SF-UniDA) addresses this challenge. Existing methods mainly rely on self-training with pseudo-labels, yet the relationship between pseudo-labeling and adaptation outcomes has not been studied yet. To bridge this gap, we conduct a systematic analysis through controlled experiments with simulated pseudo-labeling, offering valuable insights into pseudo-labeling for online SF-UniDA. Our findings reveal a substantial gap between the current state-of-the-art and the upper bound of adaptation achieved with perfect pseudo-labeling. Moreover, we show that a contrastive loss enables effective adaptation even with moderate pseudo-label accuracy, while a cross-entropy loss, though less robust to pseudo-label errors, achieves superior results when pseudo-labeling approaches perfection. Lastly, our findings indicate that pseudo-label accuracy is in general more crucial than quantity, suggesting that prioritizing fewer but high-confidence pseudo-labels is beneficial. Overall, our study highlights the critical role of pseudo-labeling in (online) SF-UniDA and provides actionable insights to drive future advancements in the field. Our code is available at https://github.com/pascalschlachter/PLAnalysis.
@@ I Introduction
    Existing SF-UniDA methods mainly rely on two different types of loss functions: contrastive loss [2, 3] and cross-entropy loss [1, 4, 5]. How do they differ, particularly in terms of robustness against erroneous pseudo-labels?
@@
    Building on the idea introduced by [6], we exclude samples with uncertain pseudo-labels from the adaptation process in our recently proposed online SF-UniDA methods [2, 3]. The underlying intuition is that pseudo-label quality (accuracy) is more important than the quantity of samples used for adaptation. But is this truly the case?
@@ II Related Work
Source-free UniDA (SF-UniDA) was first introduced by [1] with the Global and Local Clustering (GLC) method, which employs clustering-based pseudo-labeling. Adaptation is performed using a combination of cross-entropy loss and a kNN-based loss function. [4] extends this approach by incorporating an additional contrastive loss to enhance adaptation. Building on this foundation, [5] introduces Learning Decomposition (LEAD), a novel pseudo-labeling technique that separates features into source-known and -unknown components. This especially enables the use of individual rejection thresholds for unknown class identification during pseudo-labeling and allows the integration of confidence-weighting into the cross-entropy loss. Despite these variations, all these methods rely on entropy-based rejection for unknown classes during inference.
@@ III-A1 Pseudo-Label Simulation
A predefined pseudo-label quantity $q$ is then achieved by selecting the top $q\%$ of samples with the smallest entropy distances $d_{i}$.
@@ III-A2 Loss Functions
The second approach primarily relies on a cross-entropy loss for adaptation, where the unknown class is represented by uniformly distributed label vectors [1, 4, 5]. Although these works combine cross-entropy with additional loss functions, we opt to use the cross-entropy loss alone in this study, as the other losses are either unsuitable for online settings—such as a kNN-based loss—or specifically designed to support particular pseudo-labeling strategies.
@@ III-A4 Implementation details
We evaluate ODA and OPDA using the H-score, i.e. the harmonic mean of known and unknown class accuracies, and PDA with standard accuracy. Each experiment is repeated three times, and we report the average performance.
@@ IV-A Upper Bound of Adaptation (Q1)
As expected, the best results are achieved with perfect pseudo-labeling (100% pseudo-label quality and quantity). These results significantly surpass the current state-of-the-art achieved by [3], which yields an accuracy of $41.2\%$ for PDA, an H-score of $59.9\%$ for ODA, and an H-score of $60.3\%$ for OPDA. Specifically, our simulation demonstrates that perfect pseudo-labeling achieves an accuracy of 75.6% for PDA, an H-score of $74.8\%$ for ODA, and an H-score of $71.1\%$ for OPDA using the contrastive loss-based approach. Even more strikingly, the cross-entropy loss-based approach reaches $91.9\%$ for PDA, $76.7\%$ for ODA, and $82.5\%$ for OPDA under perfect pseudo-labeling conditions. Hence, these results reveal a substantial gap between the upper bound of adaptation performance and the current state-of-the-art, highlighting the significant potential for improvement in existing methods. This margin underscores the critical role of pseudo-labeling accuracy and reliability in achieving optimal adaptation outcomes, suggesting that advancements in pseudo-labeling strategies could lead to substantial performance gains across all scenarios.
@@ IV-B Pseudo-Label Requirements for Effective Adaptation (Q2)
The minimum pseudo-labeling requirements for effective adaptation depend heavily on the chosen loss function. Our results show that for the contrastive loss-based approach, adaptation remains effective even with a pseudo-label accuracy as low as $30\%$, as long as the pseudo-label quantity is sufficiently high ($\geq 40\%$). In contrast, when using cross-entropy loss, adaptation performance deteriorates much more rapidly as pseudo-label quality decreases, resulting in significantly higher minimum requirements. Specifically, a pseudo-label accuracy of at least $70\%$ is necessary to achieve improvements over the source-only model across all category shift scenarios.
@@ IV-C Comparison of Contrastive and Cross-Entropy Loss (Q3)
As already observed in the previous subsections, the contrastive loss-based approach is significantly more robust to erroneous pseudo-labels while, on the other hand, the cross-entropy loss is able to achieve substantially better results if the pseudo-label quality approaches $100\%$.
In a closer analysis, we find that the poor performance of cross-entropy loss-based adaptation with low pseudo-label quality stems from the fact that it classifies nearly all samples as unknown. This behavior is not solely caused by the pseudo-labeling process misclassifying known classes as unknown, as setting $\alpha=0$ (i.e., disallowing falsely unknown pseudo-labels) results in only small improvements. Instead, the issue appears to be inherent to the nature of the cross-entropy loss. We hypothesize that, due to its direct enforcement of desired model outputs, even small amounts of incorrect pseudo-labels can confuse the model, leading to many uncertain predictions that are subsequently interpreted as unknown.

### v2

@@ Abstract
A domain (distribution) shift between training and test data often hinders the real-world performance of deep neural networks, necessitating unsupervised domain adaptation (UDA) to bridge this gap. Online source-free UDA has emerged as a solution for practical scenarios where access to source data is restricted and target data is received as a continuous stream. However, the open-world nature of many real-world applications additionally introduces category shifts meaning that the source and target label spaces may differ. Online source-free universal domain adaptation (SF-UniDA) addresses this challenge. Existing methods mainly rely on self-training with pseudo-labels, yet the relationship between pseudo-labeling and adaptation outcomes has not been studied yet. To bridge this gap, we conduct a systematic analysis through controlled experiments with simulated pseudo-labeling, offering valuable insights into pseudo-labeling for online SF-UniDA. Our findings reveal a substantial gap between the current state-of-the-art and the upper bound of adaptation achieved with perfect pseudo-labeling. Moreover, we show that a contrastive loss enables effective adaptation even with moderate pseudo-label accuracy, while a cross-entropy (CE) loss, though less robust to pseudo-label errors, achieves superior results when pseudo-labeling approaches perfection. Lastly, our findings indicate that pseudo-label accuracy is in general more crucial than quantity, suggesting that prioritizing fewer but high-confidence pseudo-labels is beneficial. Overall, our study highlights the critical role of pseudo-labeling in (online) SF-UniDA and provides actionable insights to drive future advancements in the field. Our code is available at https://github.com/pascalschlachter/PLAnalysis.
@@ I Introduction
    Existing SF-UniDA methods primarily rely on two types of loss functions: contrastive loss [2, 3] and cross-entropy (CE) loss [1, 4, 5]. How do they differ, particularly in terms of robustness to erroneous pseudo-labels?
@@
    Following [6], we exclude samples with uncertain pseudo-labels from the adaptation in our recent online SF-UniDA methods [2, 3]. The underlying intuition is that pseudo-label quality (accuracy) is more important than the quantity of samples used for adaptation. Does empirical evidence validate this hypothesis in practice?
@@ II Related Work
Source-free UniDA (SF-UniDA) was first introduced by [1] with the Global and Local Clustering (GLC) method, which employs clustering-based pseudo-labeling. Adaptation is performed using a combination of CE loss and a kNN-based loss function. [4] extends this approach by incorporating an additional contrastive loss to enhance adaptation. Building on this foundation, [5] introduces Learning Decomposition (LEAD), a novel pseudo-labeling technique that separates features into source-known and -unknown components. This especially enables the use of individual rejection thresholds for unknown class identification during pseudo-labeling and allows the integration of confidence-weighting into the CE loss. Despite these variations, all these methods rely on entropy-based rejection for unknown classes during inference.
@@ III-A1 Pseudo-Label Simulation
A certain pseudo-label quantity $q$ is then achieved by selecting the top $q\%$ of samples with the smallest entropy distances $d_{i}$.
@@ III-A2 Loss Functions
The second approach primarily relies on a CE loss, where the unknown class is represented by uniformly distributed label vectors [1, 4, 5]. Although these works combine CE with additional loss functions, we opt to use the CE loss alone in this study, as the other losses are either unsuitable for online settings—such as a kNN-based loss—or specifically designed to support particular pseudo-labeling strategies.
@@ III-A4 Implementation details
We evaluate PDA using standard accuracy, while for ODA and OPDA, we report the H-score:
[…]
which is the harmonic mean of the known and unknown class accuracies, $acc_{k}$ and $acc_{u}$. Each experiment is repeated three times, and we report the average performance.
@@ IV-A Upper Bound of Adaptation (Q1)
As expected, the best results are achieved with perfect pseudo-labeling (100% pseudo-label quality and quantity). These results significantly surpass the current state-of-the-art achieved by [3], which yields an accuracy of $41.2\%$ for PDA, an H-score of $59.9\%$ for ODA, and an H-score of $60.3\%$ for OPDA. Specifically, our simulation demonstrates that perfect pseudo-labeling achieves an accuracy of 75.6% for PDA, an H-score of $74.8\%$ for ODA, and an H-score of $71.1\%$ for OPDA using the contrastive loss-based approach. Even more strikingly, the CE loss-based approach reaches $91.9\%$ for PDA, $76.7\%$ for ODA, and $82.5\%$ for OPDA with ideal pseudo-labeling. Hence, these results reveal a substantial gap between the upper bound of adaptation performance and the current state-of-the-art, highlighting the significant potential for improvement in existing methods. This margin underscores the critical role of pseudo-labeling accuracy and reliability in achieving optimal adaptation outcomes, suggesting that advancements in pseudo-labeling strategies could lead to substantial performance gains across all scenarios.
@@ IV-B Pseudo-Label Requirements for Effective Adaptation (Q2)
The minimum pseudo-labeling requirements for effective adaptation depend heavily on the chosen loss function. Our results show that for the contrastive loss-based approach, adaptation remains effective even with a pseudo-label accuracy as low as $30\%$, as long as the pseudo-label quantity is sufficiently high ($\geq 40\%$). In contrast, when using CE loss, adaptation performance deteriorates much more rapidly as pseudo-label quality decreases, resulting in significantly higher minimum requirements. Specifically, a pseudo-label accuracy of at least $70\%$ is necessary to achieve improvements over the source-only model across all category shift scenarios.
@@ IV-C Comparison of Contrastive and Cross-Entropy Loss (Q3)
As already observed previously, the contrastive loss-based approach is significantly more robust to erroneous pseudo-labels while the CE loss is able to achieve substantially better results if the pseudo-label quality approaches $100\%$.
In a closer analysis, we find that the poor performance of the CE loss with low pseudo-label quality stems from the fact that it classifies nearly all samples as unknown. This behavior is not solely caused by the pseudo-labeling process misclassifying known classes as unknown, as setting $\alpha=0$ (i.e., disallowing falsely unknown pseudo-labels) results in only small improvements. Instead, the issue appears to be inherent to the nature of the CE loss. We hypothesize that, due to its direct enforcement of desired model outputs, even small amounts of incorrect pseudo-labels can confuse the model, leading to many uncertain predictions that are subsequently interpreted as unknown.